OpenStudy (anonymous):

k(x)= \[\int\limits_{\cos \alpha}^{\sin ^2x^{16}} 1/(3r+r^2)\] and alpha is constant. Use the fundamental theorem of calculus to find dk/dx

OpenStudy (anonymous):

\[\frac{1}{3\sin^2x^{16}+\sin^4x^{16}} - \frac{1}{3\cos\alpha + \cos^2\alpha}\]Probably.

OpenStudy (anonymous):

no its not the one, i dont have the option in this multiple choice question

OpenStudy (anonymous):

no its not the one, i dont have this option in this multiple choice question

OpenStudy (turingtest):

Ishaan forgot to square the other x in the denom on the first term should be x^32 is that a choice ?

OpenStudy (turingtest):

... or how about is this integral even with respect to dr ? or something else?

OpenStudy (anonymous):

\[\frac{2\sin x^{16}\cos x^{16} \times 16x^{15}}{3\sin^2x^{16} + \sin^4x^{16}}\]Sorry, for the earlier mistake.

OpenStudy (anonymous):

Which can be simplified to \[\frac{32 x^{15} \cot \left(x^{16}\right)}{\sin ^2\left(x^{16}\right)+3} \]